# Multi-Group Proportional Representation

### v1

@@ 3 Computing Multi-Group Proportional Representation
Proposition 1, proved in Appendix A, bounds the deviation between the empirical MPR computed over the curated dataset $\mathcal{D}_{C}$ drawn i.i.d. from reference distribution $Q$ and the true MPR measured over $Q$.
@@
In the remainder of the paper, we assume that MPR is computed against a fixed curated dataset $\mathcal{D}_{C}$ of size $m$. We consider four specific instantiations of the set of representation statistics $\mathcal{C}$: (i) $\mathcal{C}$ is closed under scalar multiplication, i.e., if $c\in\mathcal{C}$, then $\lambda c\in\mathcal{C}$ for any $\lambda\in\mathbb{R}$, (ii) $\mathcal{C}$ is a set of functions taking values in $\{-1,1\}$; (iii) $\mathcal{C}$ consists of linear functions; and (iv) $\mathcal{C}$ consists of functions in a Reproducing Kernel Hilbert Space (RKHS). Next, we show that for cases (i) and (ii) MPR can be computed using calls to an oracle that performs regression over $\mathcal{C}$ and, for linear functions (iii), MPR has a simple closed-form expression. Due to its technical nature, we defer the calculation of MPR for functions in an RKHS to Appendix B.2.
@@
When $\mathcal{C}$ is closed under multiplication or consists of binary functions with values ${-1,1}$, MPR can be expressed as an MSE minimization. This enables us to compute MPR by leveraging existing black-box oracles that perform regression under quadratic loss, such as regressors/classifiers implemented in ML libraries like scikit-learn [68].
@@
Most classes of regression functions in $\mathbb{R}$ are closed under scalar multiplication (e.g., multiplying the output of a decision tree regressor by a constant still yields a decision tree regressor). However, in this case, it follows from (4) that MPR is unbounded. Even for bounded $\mathcal{C}$, without proper normalization MPR can scale with dataset size, which is undesirable for a representational measure. We constrain $\mathcal{C}$ to a set of normalized functions for MPR to be bounded and independent of dataset size. This constraint also allows us to cast MPR as an MSE minimization over $\mathcal{C}$. We formally state this result next, proven in Appendix A, which will be applied in Section 4 to develop a cutting-plane-based algorithm for ensuring MPR in retrieval called MAPR.
@@
Let $\mathcal{C}^{\prime}\triangleq\{c\in\mathcal{C}\mid\sum_{i=1}^{m+n}c(\boldsymbol{x}_{i})^{2}=\frac{mk}{m+k}\}$ where $\mathcal{C}$ is closed under scalar multiplication. Let $c(\mathbf{X})=[c(\mathbf{x_{1}}),...,c(\mathbf{x_{n+m}})]$. Then
$0\leq\mathrm{MPR}(\mathcal{C}^{\prime},\mathcal{R},\mathcal{D}_{C})\leq 1,$ (6)
and for any
@@
let $\hat{c}:\mathcal{X}\rightarrow\mathbb{R}$ be defined as $\hat{c}(\mathbf{x})=\frac{\sqrt{C}}{\|c^{*}(\mathbf{X})\|_{2}}c^{*}(\mathbf{x})$. Then, we have that
@@
Recall that each item in the retrieval or curated datasets is of the form $x_{i}=(e_{i},g_{i})$, where $e_{i}$ is an embedding and $g-i$ are labels. When $x_{i}\in\mathbb{R}^{l}$ (i.e., the labels have numerical values), we can consider $\mathcal{C}$ as a linear set of functions over retrieved items. In this case, MPR enjoys a closed-form expression, as stated in the next proposition.
@@
where $\mathbf{X}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\intercal}$ is the SVD of $\mathbf{X}$ and $\mathbf{U}_{l}\in\mathbb{R}^{(m+n)\times l}$ are the left singular vectors in $\mathbf{U}$ corresponding to the top-$l$ largest singular values. [^1] For analysis of functions in an RKHS, see Appendix B.2
The above proposition can be directly adapted to linear functions defined only over embeddings or a subset of features of retrieved items. The closed-form expression for MPR in (9) also allows MPR-constrained retrieval to be computed by a quadratic program, as discussed in Appendix B.1.
@@ 4 Promoting Multi-Group Proportional Representation in Retrieval Tasks
Next, we develop an optimization framework for retrieving the $k$ most similar items in a vector database to a given query $\mathbf{q}$ while satisfying a target MPR threshold of $\rho$. We formulate the retrieval goal as maximizing the average similarity between a query and retrieved items, given by $s_{i}=\kappa(\boldsymbol{x}^{r}_{i},\mathbf{q})$ (recall Section 2 for notation), where $\kappa(\boldsymbol{x}^{r}_{i},\mathbf{q})$ is given by cosine similarity in our experiments. The problem of maximizing utility in retrieval while satisfying an MPR constraint (expressed as Eq. (4)) can be formulated as the integer program:
@@
When $\mathcal{C}$ is given by normalized linear functions as in Proposition 4, the integer constraints can be relaxed to $\mathbf{a}\in[0,1]^{n}$ and the optimization can be approximated via standard quadratic solvers; see Appendix B.1.
Algorithm 1 MAPR (Multi-group Aware Proportional Retrieval)
@@
3:     $\mathbf{a}\leftarrow\arg\max_{\mathbf{a}\in[0,1]^{m}}\mathbf{a}^{\intercal}\mathbf{s}$ subject to $\left|\frac{1}{k}\sum_{i=1}^{n}a_{i}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|\leq\rho$ for $c\in\widetilde{\mathcal{C}}$
@@
Multi-Group Aware Retrieval.
We approximate (10) via the Multi-Group Aware Proportional Retrieval Algorithm (MAPR), described in Algorithm 1. MAPR is essentially a cutting-plane method that aims to find a set of $k$ items that maximize utility while approximately satisfying an MPR constraint of $\rho$. The algorithm iterates between (i) a call to an oracle that computes MPR and returns a function $c\in\mathcal{C}$ that achieves the supremum in (4) and (ii) a call to a linear program (LP) solver that approximates the top-$k$ most similar items to a query subject to linear constraints on the violation measured by $c$. Each oracle call adds an additional constraint to the LP solver which, in turn, approximates the top-$k$ items under an increasing set of constraints. The solution of the LP is rounded to satisfy the integer constraint $\mathbf{a}\in\{0,1\}^{n}$. In our implementation, we assume that MPR can be computed via MSE minimization (cf. Section 3) – in which case we consider the normalized class $\mathcal{C^{\prime}}$. The oracle call consists of running a black-box quadratic loss minimization over $\mathcal{C}$.
Interestingly, we observe that, despite relaxing the integer constraints in each LP call, the solution to the relaxed problem is very sparse and (after rounding) approximates well the solution of the ideal integer program (10) for moderate values of $\rho$. However, the method can fail to accurately approximate the IP solution when $\rho$ is small. We present a more detailed analysis of MAPR in Appendix B, which discusses potential convergence issues with the cutting plane method as well as stopping conditions.
@@ 5 Numerical Experiments
In this section, we show that MAPR is effective in promoting more proportional representation across intersectional groups while preserving similarity between retrieved items and a given query. Notably, MAPR Pareto-dominates competing benchmarks in terms of achieved MPR gap and utility.
@@
Table 1: Retrieval averaged over 10 queries with 50 retrievals on UTKFace dataset [39] for MAPR, $k$-NN, and MMR [27]. Entries indicate average percentage representation in retrieved items. MAPR is able to balance representation across classes, whereas $k$-NN and MMR miss intersectional groups (highlighted in red).
@@
MAPR  (Ours) Male 7.8 $\pm$ 3.74 10.2 $\pm$ 2.90 13.0 $\pm$ 2.56 11.8 $\pm$ 2.44 7.2 $\pm$ 3.70 50 $\pm$ 0
@@
Datasets.
We conduct retrieval over three image datasets of faces: CelebA [37], which includes labels for gender, age, and various other attributes, UTKFace [39], which contains gender, age, and race attributes, and Occupations [2], which contains gender attributes. We compute MPR using FairFace [38] as the curated dataset $\mathcal{D}_{C}$, since it is a carefully designed dataset of faces with subgroup attributes for race, gender, and age. In all cases, each dataset entry consists of an image (CLIP) embedding $\mathbf{e}_{i}$ and a set of labels $\mathbf{g}_{i}$.
Benchmarks.
We compare our method with four baselines. DebiasClip [20] modifies text queries with a prepended learned embedding for debiased retrieval but does not allow for tunable control of the representation-similarity trade-off, so it only results in a single retrieval set over which we compute similarity and MPR. ClipClip [18] trims features from CLIP embeddings that have high “mutual information” with gender, allowing for control of how many features are clipped from the embedding. PBM [29] post-processes CLIP embeddings to mitigate bias by predicting gender attributes and subsampling from each gender equally. PBM also allows tuning the representation-similarity tradeoff by controlling the likelihood of sampling at random or in a balanced manner from the retrieval dataset. Finally, the state-of-the-art for fair retrieval is the PATHS algorithm [27], which uses a set of adjective-noun keywords to capture a broad set of identity vectors in vision-language models, and then greedily samples using the MMR [45] algorithm to navigate the representation-similarity tradeoff. However, the authors’ method of computing PATHS embeddings is not public, so we used CLIP embeddings to replicate their algorithm. We refer to this method as MMR. For a detailed discussion of these algorithms and their approaches to representation, see Appendix C.
Experimental Setup.
We consider two classes of representation statistics $\mathcal{C}$: linear regression and decision trees. In both cases, we compute MPR over $\mathcal{C}^{\prime}$, i.e., $\mathcal{C}$ normalized over the retrieval and curated datasets (see Prop. 3). To accelerate retrieval, we query each of our retrieval datasets for the top 10k items according to the prompt ‘‘A photo of a {query}" in terms of raw similarity using FAISS [69]. We similarly retrieve the top 10k items from FairFace as the query-conditioned curated dataset $\mathcal{D}_{C}$. One reason we first filter for the top 10k items is the runtime of MMR, which is a greedy algorithm that traverses the full dataset for each retrieval $k$ (even for 10k entries and $k=50$ retrieved items, MMR takes hours to run for 10 queries). We aggregate results over 10 queries for occupations suggested by ChatGPT 3.5 (Appendix E).
In order to evaluate representation over interpretable population groups, we consider that representation statistics $\mathcal{C}$ are computed only over labels $\mathbf{g}_{i}$ given in FairFace. These labels are not present in all retrieval datasets. Thus, for CelebA and Occupations, we train linear probes on FairFace’s CLIP embeddings to predict their race and age attributes, and when using UTKFace, we map from FairFace’s race labels to UTKFace’s race labels by mapping Southeast Asian and East Asian to Asian and Middle Eastern and Latino_Hispanic to Other. While the fair ML community should engage in broader discussions addressing the ethics of predicting sensitive attributes from CLIP embeddings and the issues surrounding the grouping and re-grouping of diverse racial identities, we acknowledge that these are larger-scale issues that our work does not presume to address. Instead, our experiments aim to demonstrate the flexibility and scalability of the MPR metric in handling complex intersectional groups as datasets become more representative.
We retrieve $k=50$ items for each of the above queries. For a given function class and query, we compute the baseline MPR and average cosine similarity for the top 50 most similar items. Then, we conduct a parameter sweep over $\rho$ starting from this max-MPR value, inputting each value to MAPR in Algorithm 1.
Results.
We include plots for three datasets for linear regression models in Fig. 1 with additional results for decision trees and MLPs in Appendix D. We observe that MAPR Pareto-dominates benchmark methods, and is able to preserve similarity while reaching lower levels of MPR. Interestingly, methods based on directly modifying embeddings or queries for a single group attribute such as ClipClip and DebiasClip only partially reduce MPR at a high utility cost. The most competitive method to ours is MMR, which is our attempt to replicate PATHS. Though MMR is competitive to ours for large MPR, it fails to achieve small values of MPR relative to MAPR. It is also notable that MAPR can drive MPR to near zero for UTKFace, yet gaps remain in CelebA and Occupations, indicating biases in these datasets relative to FairFace.
In Fig. 3, we evaluate the quality of MAPR in terms of the precision of the retrieved images on the Occupations dataset, which has ground truth occupation labels we can query for. We observe that MAPR  retains similar precision values across levels of representation with existing methods and more fully explores the precision-representation trade-off curve. In Fig. 3, we show results for the query “A photo of a chief executive officer” with results for additional queries in Appendix D.
In the case where $\mathcal{C}$ consists of linear models, Proposition 3 provides a closed-form solution to MPR. We evaluate MAPR against a quadratic program directly optimizing for Prop 3 as demonstrated in Appendix B.1 and also compare results when rounding the relaxed problem to actually retrieve k items (Top-k). Fig 3 demonstrates that with fifty iterations, MAPR is near-optimal at all levels of representation to a direct optimization of the closed form of MPR.
In Table 1, we construct a synthetic curation set of perfectly balanced group attributes. This represents a stylized world where proportional representation is equal representation across all groups. Our results demonstrate that MAPR can exactly achieve our equal representation goals when provided with an idealized curation dataset, whereas prior baselines – see DebiasCLIP, ClipCLIP, and PBM in Appendix Table 3) – underrepresent several intersectional groups.
Figure 1: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR averaged over 10 queries for $k=50$ images retrieved. From Left-to-Right: CelebA, UTKFaces, Occupations. Values are normalized so $k$-NN achieves point (1,1) in each case. MAPR Pareto-dominates baselines and significantly closes the MPR gap.
Figure 2: Precision of MAPR on the Occupations dataset for occupation “Chief Executive Officer". Our method maintains similar precision-representation performance when compared with baselines.
Figure 3: Comparison of MAPR (iterative linear program) to quadratic program optimization of closed form MPR for linear regression. For most values of $\rho$, MAPR is near-optimal to the QP. In addition, selecting the top-k entries to conduct retrieval from the LP does not incur significant cost in terms of similarity.
6 Concluding Remarks and Limitations

### v2

@@ 3 Computing Multi-Group Proportional Representation
Proposition 1, proved in Appendix C, bounds the deviation between the empirical MPR computed over the curated dataset $\mathcal{D}_{C}$ drawn i.i.d. from reference distribution $Q$ and the true MPR measured over $Q$.
@@
In the remainder of the paper, we assume that MPR is computed against a fixed curated dataset $\mathcal{D}_{C}$ of size $m$. We consider four specific instantiations of the set of representation statistics $\mathcal{C}$: (i) $\mathcal{C}$ is closed under scalar multiplication, i.e., if $c\in\mathcal{C}$, then $\lambda c\in\mathcal{C}$ for any $\lambda\in\mathbb{R}$, (ii) $\mathcal{C}$ is a set of functions taking values in $\{-1,1\}$; (iii) $\mathcal{C}$ consists of linear functions; and (iv) $\mathcal{C}$ consists of functions in a Reproducing Kernel Hilbert Space (RKHS). Next, we show that for cases (i) and (ii) MPR can be computed using calls to an oracle that performs regression over $\mathcal{C}$ and, for linear functions (iii), MPR has a simple closed-form expression. Due to its technical nature, we defer the calculation of MPR for functions in an RKHS to Appendix D.2.
@@
When $\mathcal{C}$ is closed under multiplication or consists of binary functions with values ${-1,1}$, MPR can be expressed as an MSE minimization. This enables us to compute MPR by leveraging existing black-box oracles that perform regression under quadratic loss, such as regressors implemented in scikit-learn [69].
@@
Most classes of regression functions in $\mathbb{R}$ are closed under scalar multiplication (e.g., multiplying the output of a decision tree regressor by a constant still yields a decision tree regressor). However, in this case, it follows from (4) that MPR is unbounded. Even for bounded $\mathcal{C}$, without proper normalization MPR can scale with dataset size, which is undesirable for a representational measure. We constrain $\mathcal{C}$ to a set of normalized functions for MPR to be bounded and independent of dataset size. This constraint also allows us to cast MPR as an MSE minimization over $\mathcal{C}$. We formally state this result next, proven in Appendix C, which will be applied in Section 4 to develop a cutting-plane-based algorithm for ensuring MPR in retrieval called MOPR.
@@
Let $\mathcal{C}^{\prime}\triangleq\{c\in\mathcal{C}\mid\sum_{i=1}^{m+n}c(\boldsymbol{x}_{i})^{2}=\frac{mk}{m+k}\}$ where $\mathcal{C}$ is closed under scalar multiplication. Let $c(\mathbf{X})=[c(\mathbf{x_{1}}),...,c(\mathbf{x_{n+m}})]$. Then $0\leq\mathrm{MPR}(\mathcal{C}^{\prime},\mathcal{R},\mathcal{D}_{C})\leq 1,$ and for any
@@
let $\hat{c}:\mathcal{X}\rightarrow\mathbb{R}$ be defined as $\hat{c}(\mathbf{x})=\frac{\sqrt{mk/(m+k)}}{\|c^{*}(\mathbf{X})\|_{2}}c^{*}(\mathbf{x})$. Then, we have that
@@
Recall that each item in the retrieval or curated datasets is of the form $x_{i}=(e_{i},g_{i})$, where $e_{i}$ is an embedding and $g_{i}$ are labels. When $x_{i}\in\mathbb{R}^{l}$ (i.e., the labels have numerical values), we can consider $\mathcal{C}$ as a linear set of functions over retrieved items. In this case, MPR enjoys a closed-form expression, as stated in the next proposition.
@@
where $\mathbf{X}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\intercal}$ is the SVD of $\mathbf{X}$ and $\mathbf{U}_{l}\in\mathbb{R}^{(m+n)\times l}$ are the left singular vectors in $\mathbf{U}$ corresponding to the top-$l$ largest singular values. [^1] For analysis of functions in an RKHS, see Appendix D.2
Proposition 4 can be directly adapted to linear functions defined only over embeddings or a subset of features of retrieved items. The closed-form expression for MPR in (8) also allows MPR-constrained retrieval to be computed by a quadratic program, as discussed in Appendix D.1.
@@ 4 Promoting Multi-Group Proportional Representation in Retrieval Tasks
We develop an optimization framework for retrieving the $k$ most similar items in a vector database to a given query $\mathbf{q}$ while satisfying a target MPR threshold of $\rho$. We formulate the retrieval goal as maximizing the average similarity between a query and retrieved items, given by $s_{i}=\kappa(\boldsymbol{x}^{r}_{i},\mathbf{q})$ (recall Section 2 for notation), where $\kappa(\boldsymbol{x}^{r}_{i},\mathbf{q})$ is given by cosine similarity in our experiments. The problem of maximizing utility in retrieval while satisfying an MPR constraint (expressed as Eq. (4)) can be formulated as the integer program:
@@
When $\mathcal{C}$ is given by normalized linear functions as in Proposition 4, the integer constraints can be relaxed to $\mathbf{a}\in[0,1]^{n}$ and the optimization can be approximated via standard quadratic solvers; see Appendix D.1.
Algorithm 1 MOPR (Multi-group Optimized Proportional Retrieval)
@@
3:    $\mathbf{a}\leftarrow\arg\max_{\mathbf{a}\in[0,1]^{n}}\mathbf{a}^{\intercal}\mathbf{s}$ subject to $\left|\frac{1}{k}\sum_{i=1}^{n}a_{i}c(\boldsymbol{x}^{r}_{i})-\frac{1}{m}\sum_{i=1}^{m}c(\boldsymbol{x}^{c}_{i})\right|\leq\rho$ for $c\in\widetilde{\mathcal{C}}$
@@
Multi-Group Optimized Proportional Retrieval.
We approximate (9) via the Multi-Group Optimized Proportional Retrieval Algorithm (MOPR), described in Algorithm 1. MOPR is essentially a cutting-plane method that aims to find a set of $k$ items that maximize utility while approximately satisfying an MPR constraint of $\rho$. The algorithm iterates between (i) a call to an oracle that computes MPR and returns a function $c\in\mathcal{C}$ that achieves the supremum in (4) and (ii) a call to a linear program (LP) solver that approximates the top-$k$ most similar items to a query subject to linear constraints on the violation measured by $c$. Each oracle call adds an additional constraint to the LP solver which, in turn, approximates the top-$k$ items under an increasing set of constraints. The solution of the LP is rounded to satisfy the integer constraint $\mathbf{a}\in\{0,1\}^{n}$. In our implementation, we assume that MPR can be computed via MSE minimization (cf. Section 3) – in which case we consider the normalized class $\mathcal{C^{\prime}}$. The oracle call consists of running a black-box quadratic loss minimization over $\mathcal{C}$ and normalizing.
Interestingly, we observe that, despite relaxing the integer constraints in each LP call, the solution to the relaxed problem is very sparse and (after rounding) approximates well the solution of the ideal integer program (9) for moderate values of $\rho$. However, the method can fail to accurately approximate the IP solution when $\rho$ is small. We present a more detailed analysis of MOPR in Appendix D, which discusses potential convergence issues with the cutting plane method as well as stopping conditions.
@@ 5 Numerical Experiments
In this section, we show that MOPR is effective in promoting more proportional representation across intersectional groups while preserving similarity between retrieved items and a given query. Notably, MOPR Pareto-dominates competing benchmarks in terms of achieved MPR gap and utility.
Figure 1: Fraction of Top-$k$ cosine similarity vs Fraction of Top-$k$ MPR averaged over 10 queries for $k=50$ images retrieved. From Left-to-Right: CelebA, UTKFaces, Occupations. Values are normalized so Top-$k$ MPR and similarity is the point (1,1). MOPR Pareto-dominates baselines and significantly closes the MPR gap.
Datasets.
We conduct retrieval over three image datasets of faces: CelebA [38], which includes labels for gender, age, and various other attributes, UTKFace [40], which contains gender, age, and race attributes, and Occupations [2], which contains gender attributes. We compute MPR for equal representation by constructing a synthetic dataset $\mathcal{D}_{C}$ balanced over all attributes and for proportional representation using FairFace [39] as the curated dataset $\mathcal{D}_{C}$, since it is a carefully designed dataset of faces with subgroup attributes for race, gender, and age. In all cases, each dataset entry consists of an image (CLIP) embedding $\mathbf{e}_{i}$ and a set of labels $\mathbf{g}_{i}$.
Benchmarks.
We compare our method with four baselines. DebiasCLIP [20] modifies text queries with a prepended learned embedding for debiased retrieval but does not allow for tunable control of the representation-similarity trade-off, so it only results in a single retrieval set over which we compute similarity and MPR. CLIP-Clip [18] trims features from CLIP embeddings that have high “mutual information” with gender, allowing for control of how many features are clipped from the embedding. PBM [29] post-processes CLIP embeddings to mitigate bias by predicting gender attributes and subsampling from each gender equally. PBM also allows tuning the representation-similarity tradeoff by controlling the likelihood of sampling at random or in a balanced manner from the retrieval dataset. Finally, the state-of-the-art for fair retrieval first constructs PATHS embeddings [27], which use a set of adjective-noun keywords to capture a broad set of identity vectors in vision-language models, and then greedily samples using the MMR [45] algorithm to navigate the representation-similarity tradeoff. However, the authors’ method of computing PATHS embeddings is not public, so we used CLIP embeddings to replicate their algorithm. We refer to this method as MMR. For a detailed discussion of these algorithms and their approaches to representation, see Appendix E.
Experimental Setup.
We consider three classes of representation statistics $\mathcal{C}$: linear regression, decision trees, and MLPs (the last two are presented in Appendix G). In both cases, we compute MPR over $\mathcal{C}^{\prime}$, i.e., $\mathcal{C}$ normalized over the retrieval and curated datasets (see Prop. 3). We report results over 10 queries for occupations suggested by ChatGPT 3.5 (Appendix H) and an additional 45 occupations for the the Occupations dataset. To accelerate retrieval, we query each of our retrieval datasets for the top 10k items according to the prompt ‘‘A photo of a {query}" in terms of raw similarity using FAISS [70]. We similarly retrieve the top 10k items from FairFace as the query-conditioned curated dataset $\mathcal{D}_{C}$. One reason we first filter for the top 10k items is the runtime of MMR, which is a greedy algorithm that traverses the full dataset for each retrieval $k$ (even for 10k entries and $k=50$ retrieved items, MMR takes hours to run for 10 queries).
In order to evaluate representation over interpretable population groups, we consider that representation statistics $\mathcal{C}$ are computed only over labels $\mathbf{g}_{i}$ given in FairFace. These labels are not present in all retrieval datasets. Thus, for CelebA and Occupations, we train linear probes on FairFace’s CLIP embeddings to predict their race and age attributes, and when using UTKFace, we map from FairFace’s race labels to UTKFace’s race labels by mapping Southeast Asian and East Asian to Asian and Middle Eastern and Latino_Hispanic to Other. While the fair ML community should engage in broader discussions addressing the ethics of predicting sensitive attributes from CLIP embeddings [71] (especially given that CLIP itself has been found to be biased) and the issues surrounding the grouping and re-grouping of diverse racial identities, we acknowledge that these are larger-scale issues that our work does not presume to address.
We retrieve $k=50$ items for each of the above queries. For a given function class and query, we compute the baseline MPR and average cosine similarity given by the top 50 most similar items. Then, we conduct a parameter sweep over $\rho$ starting from this max-MPR value, inputting each value to MOPR in Algorithm 1. We normalize our results such that the Top-$k$ MPR and similarity are the point (1,1) on each graph, and each point measures the fraction of Top-$k$ MPR and similarity.
Results.
In Fig. 1 we report results for proportional representation for the query ‘‘A photo of a programmer’’ with respect to FairFace for linear regression models, with additional results for other queries, as well as decision trees and MLPs in Appendix G (Figs. 8-19). We observe that MOPR Pareto-dominates benchmark methods, and is able to preserve similarity while reaching lower levels of MPR. Methods based on directly modifying embeddings or queries for a single group attribute such as CLIP-Clip and DebiasCLIP only partially reduce MPR at a high utility cost. This is because these methods do not modify the retrieval algorithm but use an unbiased embedding and hope by chance that the retrieved items will be representational. The most competitive method to ours is the retrieval algorithm MMR, which is our attempt to replicate PATHS. Though MMR is competitive to ours for large MPR, it fails to achieve small values of MPR relative to MOPR. It is also notable that MOPR can drive MPR to near zero for UTKFace, yet gaps remain in CelebA and Occupations, indicating this gap may be caused by the use of probes to estimate group attribute labels.
In Table 1, we construct a synthetic curation set with an equal number of each group attribute. This allows us to evaluate the performance of MOPR in achieving equal representation. We conduct a similar hyperparameter sweep as above, and take the retrieval set with the minimum MPR averaged over all 10 queries to fairly compare the best possible representation of each method. [^2] Aggregate results should be handled with care. As each query-conditioned retrieval setting results in a different distribution, average-case performance does not preclude the existence of a query where baseline methods achieve superior results. See Limitations (Appendix B) for further discussion. Our results demonstrate that MOPR can exactly achieve our equal representation goals when provided with an equally-balanced curation dataset, whereas prior baselines (see DebiasCLIP, CLIP-Clip, and PBM in Appendix Table 5) underrepresent several intersectional groups.
@@ Results.
Table 1: Retrieval averaged over 10 queries with 50 retrievals on UTKFace dataset [40] for MOPR, $k$-NN, and MMR [27]. Entries indicate average percentage representation in retrieved items. MOPR is able to balance representation across classes, whereas $k$-NN and MMR miss intersectional groups (highlighted in red).
@@
MOPR  (Ours) Male 7.8 $\pm$ 3.74 10.2 $\pm$ 2.90 13.0 $\pm$ 2.56 11.8 $\pm$ 2.44 7.2 $\pm$ 3.70 50 $\pm$ 0
@@
6 Concluding Remarks